OpenStudy (anonymous):

suppose U=[1,2,3,4,5,6,7,8] is the universal set and P=[1,2,3,4]. what is P'? a.[5,6,7,8] b.[1,2,3,4,5,6,7,8] c.[1.,2,3,4] d.cannot determined your goal is to save atleast $360.00 over the next 6 weeks. How much money must you save each week in order to meet that ggoal? write and solve an inequality a.6+x ≥ 360; ≥ 354 b.60x ≤ 360; x ≤ 10 c.x/6 ≤ 360; x ≤ 2,160 d.6x ≥ 360; x ≥ 60 of 158 consumers polled some like only action movies some prefer just horror movies and some like both. if 43 people like only action, and 62 like both, how many people perfer only horror movies a.5

OpenStudy (anonymous):

suppose noodle soup is on sale for $0.75 a caan, and you have a coupon for $0.50 off your total purchase, Write a finction rule for the cost of n cans of soup. a. C(n)=0.75-0.5 b.C(n)=0.5n-0.75 c. C(n)=(0.75+n)-0.5 d. C(n)=(0.5n+n)-0.75 -19,-11,-3,5 tell whether the sequence is arithmetic. If it is, what is the common difference? a.yes; 5 b.yes; 6 c.yes; 8 d.no